ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

7. Lie Groups and Lie Algebras

1991 MS Classification: 22 Koranyi, Adam, Lehman College, Bronx, USA Equivariant differential operators on homogeneous vector bundles over G/MAN Using the standard notation for a semisimple Lie group G and a maximal parabolic subgroup MAN we consider pairs of bundles given by irreducible representations of M and characters of A. WE give an explicit description of all G-equivariant first order differential operators from sections of the first bundle to sections of the second. The results are joint work with H. M. Reimann. Using different methods, similar results have been obtained simultaneously and independently by B. Orsted.